OpenStudy (anonymous):

Find the derivative of f(x) = 5 divided by x at x = -1.

OpenStudy (anonymous):

\[f(x)= \frac{ 5 }{ x }\] \[f'(x) = \frac {-5} {x^2}\] \[f'(-1) = \frac{-5}{(-1)^2} = \frac{-5}{1} = -5 \]

OpenStudy (rock_mit182):